OpenStudy (anonymous):

The minimum and max temp. On a cold day in lollipop town can be modeled by the equation below: 2|x-6|+14=38 What is the minimum and max temperatures for this day?

OpenStudy (muzzack):

x = 18

OpenStudy (muzzack):

and x = -20

OpenStudy (nurali):

2|x − 6| + 14 = 38 Subtract 14 from both sides, 2|x - 6| + 14 - 14 = 38 - 14 2|x - 6| = 24 Divide both sides by 2, [ 2|x - 6| ] / 2 = 24 / 2 |x - 6| = 12 -12 = x - 6 = 12 -12 = x - 6 -12 + 6 = x -6 = x x - 6 = 12 x = 12 + 6 x = 18 x = -6, 18 The minimum T = -6 The maximum T = 18 [ x = −6, x = 18 ]

OpenStudy (anonymous):

thanks :)

OpenStudy (nurali):

My Pleasure.